the product o two numbers is 13.if one of the numbers is 1/2 find the other
Found 3 solutions by ewatrrr, stanbon, advanced_Learner:
Answer by ewatrrr(24785)

Previously Posted
n the other number
(1/2)n = 13
n = 26 is the other number

Answer by stanbon(75887)

the product o two numbers is 13.if one of the numbers is 1/2 find the other
-----
Let the missing number be "x"::
Equation:
(1/2)*x = 13
-----
x = 26
------------
Cheers,
Stan H.
---------

Answer by advanced_Learner(501)

xy=13
y=1/2
x*(1/2)=13
x=26